# The evolution of the galaxy stellar mass function and star formation rates in the COLIBRE simulations from redshift 17 to 0

### v1

@@ 3.4 Cosmic star formation rate density
Figure 6: Cosmic star formation rate density (CSFRD) as a function of redshift in the colibre L200m6 simulation, split into 11-dex stellar mass bins (left panel) and 11-dex halo mass bins (right panel). Differently coloured solid lines represent the contributions from each bin, except for the lowest mass bin, which is shown as a grey dotted line to highlight resolution limitations. The lower (upper) edge of the lowest (highest) stellar and halo mass bins is unbounded to include all remaining objects in the simulation at the low-mass (high-mass) end. The thick dashed orange line represents the CSFRD calculated by summing the SFRs of all subhaloes within the simulation volume. To guide the eye, the black symbols show the same observational measurements as in Fig. 5. At $0<z<3$, subhaloes with masses of $M_{\rm halo}\sim 10^{12}~\mathrm{M_{\odot}}$, which host galaxies with stellar masses of $\sim 10^{10}~\mathrm{M_{\odot}}$, dominate the CSFRD. At higher redshifts ($z>3$), the primary contribution to the CSFRD shifts progressively towards subhaloes of lower mass, while the most massive objects ($M_{\rm halo}\geq 10^{13.5}~\mathrm{M_{\odot}}$) contribute negligibly at all $z$.
@@
Finally, we note that although the CSFRD from $M_{*}\geq 10^{7}~\mathrm{M_{\odot}}$ subhaloes predicted by the m5 model becomes negligible above $z=12$ due to the lack of $M_{*}\geq 10^{7}~\mathrm{M_{\odot}}$ objects in the $50^{3}$cMpc3 volume at these very high redshifts, its total CSFRD (indicated by the solid light-blue curve) at $z=12$ remains at $\sim 10^{-3}~\mathrm{M_{\odot}\,yr^{-1}\,cMpc^{-3}}$, decreasing by an order of magnitude by $z=18$. At $8\lesssim z\lesssim 14$, this total CSFRD exceeds all JWST measurements shown in Fig. 5 by $\approx 0.5$ to $\approx 1.5$ dex, which are based on integrating the UV luminosity function down to $M_{\rm UV}=-17$. This result implies that, due to the contribution from objects too faint to be currently observed – with $M_{\rm UV}\gtrsim-17$ or, equivalently, $M_{*}\lesssim 10^{7}~\mathrm{M_{\odot}}$ – the true CSFRD in the Universe at high redshifts may be significantly higher than current observational estimates suggest.
@@
In terms of stellar mass, the CSFRD at $z\gtrsim 6$ is dominated by the lowest stellar mass bin, reflecting the resolution limit of the m6 simulation. The galaxies in this bin either contain only a few stellar particles (with a mass of $m_{\rm*}\sim 10^{6}~\mathrm{M_{\odot}}$) or have not yet formed any stars. Focussing again on the right panel, we find that the majority of these galaxies must reside in subhaloes with masses of $10^{8.5}\lesssim M_{\rm halo}/\mathrm{M_{\odot}}\lesssim 10^{9.5}$, as these subhaloes dominate the $z\gtrsim 6$ CSFRD [^9] The spike at $z=6.5$ in the CSFRD from the $10^{12.5}\leq M_{\rm halo}/\mathrm{M_{\odot}}<10^{13.5}$ bin is caused by a single starburst galaxy with $M_{\rm halo}\approx 10^{12.56}~\mathrm{M_{\odot}}$ whose SFR drops from $\approx 5\times 10^{2}~\mathrm{M_{\odot}\,yr^{-1}}$ at $z=6.5$ to $\approx 10^{2}~\mathrm{M_{\odot}\,yr^{-1}}$ by $z=6$..
@@
Figure 7: Star-forming main sequence (SFMS) in the L200m6 simulation at $z=0.1$, $0.5$, 11, 22, 44, and $6$ (different panels). The colour scale shows the fraction of galaxies per pixel with a given $M_{*}$ and SFR, normalized by the total number of galaxies across all pixels at the same $M_{*}$. The solid orange lines represent the colibre SFMS, defined as the median SFR of star-forming galaxies ($\mathrm{SFR}>0.1\,\times$ the SFR at the SFMS determined iteratively). The short-dashed light-green and red dash-dotted curves show alternative SFMS definitions based on the mode and mean of the SFR distribution at fixed $M_{*}$, respectively. For reference, the diagonal thin dotted line shows $\mathrm{sSFR}=0.2/t_{\rm H}(z)$, which is commonly used as an sSFR threshold for defining star-forming galaxies. Furthermore, also for reference, white diamonds and triangles indicate the observed SFMSs from Leja et al. (2022) and Popesso et al. (2023), respectively, while the white dashed contours mark the regions enclosing 95 per cent of the sample of star-forming galaxies from Euclid Collaboration et al. (2025). All three definitions of the SFMS in the simulation yield very similar results at all redshifts.
@@ 3.5 Evolution of the star-forming main sequence
Fig. 7 also shows the $0.2<z<3$ measurements from the Euclid Q1 data release (Euclid Collaboration et al., 2025), which is based on the Euclid Deep Fields (an effective area of $\approx 63$ deg2) and includes over $4\times 10^{6}$ objects above the 95 per cent stellar mass completeness limit up to $z=3$. Specifically, from this sample, we show the distribution of star-forming galaxies in the SFR – $M_{*}$ plane in three redshift bins (median values of $z=0.37$, $1.15$, and $1.83$), indicated by white dashed contours enclosing 95 per cent of the star-forming galaxies (i.e. $2\sigma$ of the star-forming distribution). The distribution of galaxies in the L200m6 simulation (i.e. pixels coloured in different shades of blue) largely covers the $2\sigma$ contours from Euclid Collaboration et al. (2025) at all three redshifts. Finally, Fig. 7 includes the observed SFMSs from Leja et al. (2022) and Popesso et al. (2023), indicated by white diamonds and triangles, respectively. These two SFMSs are shown for reference only, and we defer discussion of how well colibre reproduces these two (and other) observed relations to the next figure.
@@
The SFMS in the m5, m6, and m7 colibre models qualitatively matches the observational data at all examined redshifts. At $z\geq 5$, colibre reproduces both the normalization and slope of the SFMS reported by Thorne et al. (2021), Popesso et al. (2023), and Cole et al. (2025). For $1<z<4$, the colibre SFMS is $\approx 0.2-0.6$ dex below all observed SFMSs at the high-mass end ($M_{*}\gtrsim 10^{10.5}~\mathrm{M_{\odot}}$), while still broadly matching the SFMS of Leja et al. (2022) and, to a somewhat lesser degree, Koprowski et al. (2024) at $M_{*}\lesssim 10^{10.5}~\mathrm{M_{\odot}}$. Below $z=1$, the colibre SFMS becomes once again broadly consistent with Thorne et al. (2021) and Popesso et al. (2023). While colibre shows excellent convergence with resolution at $z\gtrsim 5$, at lower redshifts the higher-resolution models tend to predict lower SFRs at fixed stellar mass, leading to larger deviations from the observations.
The fact that colibre falls short of the SFMS from Thorne et al. (2021) and Popesso et al. (2023) by 0.20.60.2-0.6 dex near cosmic noon warrants further discussion. As previously mentioned, a discrepancy of this magnitude between theoretical models and observations at $1<z<4$ is common in galaxy formation studies. It has been shown that one way to mitigate these discrepancies is to mock-observe the simulated galaxies rather than relying on the internal SFRs and stellar masses predicted by the simulations. This approach involves generating realistic SEDs based on the distribution of stellar particle ages and metallicities in the simulation, which are then processed through the same observational pipelines used for real galaxies to infer SFRs and stellar masses (e.g. Katsianis et al., 2020; Nelson et al., 2021).
@@
Figure 9: Evolution of the galaxy quenched fraction from $z=0.1$ to $z=7$ (different panels). Thick solid lines show the quenched fraction in the simulations, where galaxies are considered quenched if their $\mathrm{sSFR}<0.2/t_{\rm H}(z)$. Shaded regions indicate the uncertainty in the predicted quenched fractions, calculated as the Clopper–Pearson interval at the 68 per cent confidence level. At low stellar masses ($M_{*}\lesssim 10^{9}~\mathrm{M_{\odot}}$), the solid lines become dotted to indicate that the increase in quenched fraction is likely an artefact of limited numerical resolution. The lines also become dotted at high $M_{*}$ where the number of objects per bin drops below 10. Black symbols represent a compilation of observational data from Muzzin et al. (2013), Bauer et al. (2013), Moster et al. (2018), and Behroozi et al. (2019). Overall, colibre shows good agreement with the data at all redshifts. However, at $z\sim 0$, the most massive galaxies ($M_{*}\gtrsim 10^{11.5}~\mathrm{M_{\odot}}$) appear more star-forming than observed.
@@ 3.6 Evolution of the galaxy quenched fraction
We show the fraction of quenched galaxies as a function of stellar mass in Fig. 9. Each panel corresponds to a different redshift, ranging from $z=0.1$ (top left) to $z=7$ (bottom right). The quenched fraction in the simulations (shown as solid lines) is defined as the fraction of galaxies within a given stellar mass bin with sSFRs below $0.2/t_{\rm H}(z)$ (black dotted lines in Fig. 7). Such thresholds – which compare galaxy sSFRs to a fraction of either the Hubble time or the age of the Universe at a given redshift – are commonly used in the literature to distinguish between quiescent and star-forming galaxy populations (e.g. Franx et al., 2008; Gallazzi et al., 2014; Baker et al., 2025b; Carnall et al., 2024; Russell et al., 2024). At low stellar masses ($M_{*}\lesssim 10^{9}~\mathrm{M_{\odot}}$), the solid lines become dotted to emphasize that the rise in quenched fractions may be driven by limited numerical resolution (see below). The solid lines also become dotted at the high-mass end, indicating that there are fewer than 10 objects per stellar mass bin. The shaded regions indicate the uncertainty in the quenched fractions, computed as the Clopper–Pearson interval at the 68 per cent confidence level.
@@
To isolate the role of AGN feedback, we re-ran the colibre m7 simulation in a ($50~\mathrm{cMpc}$)3 cosmological volume without SMBHs. This test confirmed that the rise in quenched fraction from $M_{*}\sim 10^{10}$ to $\sim 10^{11.5}~\mathrm{M_{\odot}}$ is driven by AGN feedback in massive haloes, while the decrease at $M_{*}\gtrsim 10^{11.5}~\mathrm{M_{\odot}}$ is likely an indication of cooling flows in the BCGs, which the colibre AGN feedback model cannot fully counteract. However, the dip at $M_{*}\gtrsim 10^{11.5}~\mathrm{M_{\odot}}$ disappears when we increase the assumed lognormal scatter on the stellar mass from $\approx 0.1$ dex (as given by equation 1 at $z\approx 0$) to $\gtrsim 0.3$ dex (see Fig. 19 in Appendix D). Similarly, the dip also disappears if the SFR is computed within an aperture of $\lesssim 10$ pkpc (comparable to the physical size at $z=0.1$ of a fibre with a diameter of a few arcseconds in surveys such as SDSS or GAMA), rather than the default $50$ pkpc aperture used for all quantities in this work (not shown).
@@
At higher redshifts ($z\gtrsim 1$), colibre remains in quantitative agreement with the observational data for $M_{*}\gtrsim 10^{9}~\mathrm{M_{\odot}}$, with the quenched fraction steadily decreasing with increasing redshift at all stellar masses. However, caution should be taken when drawing more detailed conclusions, as differences in the criteria for defining quiescent galaxy populations between the observations and colibre may influence the results. For galaxies with $M_{*}\sim 10^{9}-10^{10}~\mathrm{M_{\odot}}$, colibre predicts a nearly zero quenched fraction by $z=1$, which remains negligible also at higher redshifts. For $M_{*}\sim 10^{11}~\mathrm{M_{\odot}}$, the quenched fraction decreases from $\approx 70-80$ per cent at $z=0.1$ to 204020-40 per cent by $z=2$, and further drops to 5305-30 per cent by $z=5$.
Although the $z=0.1$ quenched fractions show good convergence with resolution for $10^{10}\lesssim M_{*}/\mathrm{M_{\odot}}\lesssim 10^{11}$, at $z\geq 2$ the m6 simulation predicts systematically lower quenched fractions (i.e. weaker quenching) than m7 in this mass range [^11] The volume of the m5 simulation does not contain a sufficient number of massive quenched galaxies at high redshifts to determine whether quenched fractions also differ systematically between m6 and m5 resolutions.. The explanation for this discrepancy is likely twofold. First, the gas surrounding SMBHs in the simulation at m6 resolution is, on average, denser than in the m7 simulation, which may increase radiative cooling losses, thereby reducing the effectiveness of AGN feedback on the host galaxy. Second, according to the colibre prescription for thermal AGN feedback, BHs inject energy into neighbouring gas particles such that the gas temperature increases by $\Delta T_{\rm AGN}$ upon injection. Because the gas particle mass is a factor of eight lower in the m6 simulation compared to m7, the energy required to heat a gas particle by the same $\Delta T_{\rm AGN}$ is also eight times lower in m6 than in m7. In other words, less energy is injected per AGN event in higher-resolution simulations, making the AGN feedback less bursty and potentially reducing its efficiency in expelling gas and quenching the galaxy. This may also explain why, at $z=0.1$, the dip in the galaxy quenching fraction at $M_{*}\gtrsim 10^{11.5}~\mathrm{M_{\odot}}$ is slightly more pronounced at m6 resolution than at m7.
@@
Satellites in colibre exhibit systematically higher quenched fractions than centrals at fixed stellar mass, a trend consistent at all resolutions. For galaxies with $M_{*}\sim 10^{9}-10^{10}~\mathrm{M_{\odot}}$, the quenched fraction of centrals ranges between 0 and $10$ per cent, whereas satellites show values of $\approx 15-40$ per cent, reflecting the important role of environmental effects. The impact of environmental effects diminishes for more massive satellites ($M_{*}\gtrsim 10^{10}~\mathrm{M_{\odot}}$), as their deeper gravitational potential wells enable them to better retain their gas reservoirs in the host’s environment. This leads to quenched fractions for these satellites becoming more similar to those of centrals at comparable stellar masses. A detailed study of environmental effects on galaxy properties will be presented in Bahé et al. (in preparation).
For $M_{*}>10^{9}~\mathrm{M_{\odot}}$, the quenched fraction for centrals shows excellent convergence across all colibre resolutions, whereas for satellites, the quenched fraction for $10^{9}\lesssim M_{*}/\mathrm{M_{\odot}}\lesssim 10^{10}$ in the m5 model is a factor of $\approx 2$ lower than in the m6 and m7 counterparts. The cause for this discrepancy is the relatively small cosmological volume used in the m5 simulation, which biases the predicted satellite quenched fraction towards lower values due to the lack of massive haloes ($M_{\rm halo}\gtrsim 10^{13.5}~\mathrm{M_{\odot}}$). In Fig. 16 of Appendix A, we show that the bias is significantly reduced when all three resolutions are compared within the same cosmological volume $(25^{3}~\mathrm{cMpc}^{3}$).

### v2

@@ 3.4 Cosmic star formation rate density
Figure 6: Cosmic star formation rate density (CSFRD) as a function of redshift in the colibre L200m6 simulation, split into 11-dex stellar mass bins (left panel) and 11-dex halo mass bins (right panel). Differently coloured solid lines represent the contributions from each bin, except for the lowest mass bin, which is shown as a grey dotted line to highlight resolution limitations. The lower (upper) edge of the lowest (highest) stellar and halo mass bins is unbounded to include all remaining objects in the simulation at the low-mass (high-mass) end. The halo mass is the Bryan & Norman (1998) mass for central subhaloes and the total gravitationally bound mass for satellites. Subhalo SFRs are obtained by summing the instantaneous SFRs of gas particles that are bound to the subhalo and within a 50 pkpc 3D spherical aperture. The thick dashed orange line represents the CSFRD calculated by summing the SFRs of all subhaloes within the simulation volume. To guide the eye, the black symbols show the same observational measurements as in Fig. 5. At $0<z<3$, subhaloes with masses of $M_{\rm halo}\sim 10^{12}~\mathrm{M_{\odot}}$, which host galaxies with stellar masses of $\sim 10^{10}~\mathrm{M_{\odot}}$, dominate the CSFRD. At higher redshifts ($z>3$), the primary contribution to the CSFRD shifts progressively towards subhaloes of lower mass, while the most massive objects ($M_{\rm halo}\geq 10^{13.5}~\mathrm{M_{\odot}}$) contribute negligibly at all $z$.
@@
Finally, we note that although the CSFRD from $M_{*}\geq 10^{7}~\mathrm{M_{\odot}}$ subhaloes predicted by the m5 model becomes negligible above $z=14$ due to the lack of $M_{*}\geq 10^{7}~\mathrm{M_{\odot}}$ objects in the $100^{3}$ cMpc3 volume at these very high redshifts, its total CSFRD (indicated by the solid light-blue curve) at $z=14$ remains at $\sim 10^{-3}~\mathrm{M_{\odot}\,yr^{-1}\,cMpc^{-3}}$, decreasing by an order of magnitude by $z=18$. At $8\lesssim z\lesssim 14$, this total CSFRD exceeds all JWST measurements shown in Fig. 5 by $\approx 0.5$ to $\approx 1.5$ dex, which are based on integrating the UV luminosity function down to $M_{\rm UV}=-17$. This result implies that, due to the contribution from objects too faint to be currently observed – with $M_{\rm UV}\gtrsim-17$ or, equivalently, $M_{*}\lesssim 10^{7}~\mathrm{M_{\odot}}$ – the true CSFRD in the Universe at high redshifts may be significantly higher than current observational estimates suggest.
@@
In terms of stellar mass, the CSFRD at $z\gtrsim 6$ is dominated by the lowest stellar mass bin, reflecting the resolution limit of the m6 simulation. The galaxies in this bin either contain only a few stellar particles (with a mass of $m_{\rm*}\sim 10^{6}~\mathrm{M_{\odot}}$) or have not yet formed any stars. Focussing again on the right panel, we find that the majority of these galaxies must reside in subhaloes with masses of $10^{8.5}\lesssim M_{\rm halo}/\mathrm{M_{\odot}}\lesssim 10^{9.5}$, as these subhaloes dominate the $z\gtrsim 6$ CSFRD [^10] The spike at $6.5<z<6.75$ in the CSFRD from the $10^{12.5}\leq M_{\rm halo}/\mathrm{M_{\odot}}<10^{13.5}$ bin is caused by a single starburst galaxy with $M_{\rm halo}\approx 10^{12.52}~\mathrm{M_{\odot}}$ ($\approx 10^{12.56}~\mathrm{M_{\odot}}$) at $z=6.75$ ($z=6.5$) whose SFR drops from $\approx 5\times 10^{2}~\mathrm{M_{\odot}\,yr^{-1}}$ at $6.5<z<6.75$ to $\approx 10^{2}~\mathrm{M_{\odot}\,yr^{-1}}$ by $z=6$..
@@
Figure 7: Star-forming main sequence (SFMS) in the L200m6 simulation at $z=0.1$, $0.5$, 11, 22, 44, and $6$ (different panels). The colour scale shows the fraction of galaxies per pixel with a given $M_{*}$ and SFR, normalized by the total number of galaxies across all pixels at the same $M_{*}$. The solid orange lines represent the colibre SFMS, defined as the median SFR of star-forming galaxies ($\mathrm{SFR}>0.1\,\times$ the SFR at the SFMS determined iteratively). The short-dashed light-green and red dash-dotted curves show alternative SFMS definitions based on the mode and mean of the SFR distribution at fixed $M_{*}$, respectively. For reference, the diagonal thin dotted line shows $\mathrm{sSFR}=0.2/t_{\rm H}(z)$, which is commonly used as an sSFR threshold for defining star-forming galaxies. Furthermore, also for reference, white diamonds and triangles indicate the observed SFMSs from Leja et al. (2022) and Popesso et al. (2023), respectively, while the white dashed contours mark the regions enclosing 95 per cent of the sample of star-forming galaxies from Euclid Collaboration et al. (2025a). All three definitions of the SFMS in the simulation yield very similar results at all redshifts.
@@ 3.5 Evolution of the star-forming main sequence
Fig. 7 also shows the $0.2<z<3$ measurements from the Euclid Q1 data release (Euclid Collaboration et al., 2025a), which is based on the Euclid Deep Fields (an effective area of $\approx 63$ deg2) and includes over $4\times 10^{6}$ objects above the 95 per cent stellar mass completeness limit up to $z=3$. Specifically, from this sample, we show the distribution of star-forming galaxies in the SFR – $M_{*}$ plane in three redshift bins (median values of $z=0.37$, $1.15$, and $1.83$), indicated by white dashed contours enclosing 95 per cent of the star-forming galaxies (i.e. $2\sigma$ of the star-forming distribution). The distribution of galaxies in the L200m6 simulation (i.e. pixels coloured in different shades of blue) largely covers the $2\sigma$ contours from Euclid Collaboration et al. (2025a) at all three redshifts. Finally, Fig. 7 includes the observed SFMSs from Leja et al. (2022) and Popesso et al. (2023), indicated by white diamonds and triangles, respectively. These two SFMSs are shown for reference only, and we defer discussion of how well colibre reproduces these two (and other) observed relations to the next figure.
@@
The SFMS in the m5, m6, and m7 colibre models qualitatively matches the observational data at all examined redshifts. At $z\geq 5$, colibre reproduces both the normalization and slope of the SFMS reported by Thorne et al. (2021), Popesso et al. (2023), and Cole et al. (2025). For $1<z<4$, the colibre SFMS is $\approx 0.2-0.6$ dex below all observed SFMSs at the high-mass end ($M_{*}\gtrsim 10^{10.5}~\mathrm{M_{\odot}}$), while still broadly matching the SFMS of Leja et al. (2022) and, to a somewhat lesser degree, Koprowski et al. (2024) at $M_{*}\lesssim 10^{10.5}~\mathrm{M_{\odot}}$. Below $z=1$, the colibre SFMS once again becomes broadly consistent with Thorne et al. (2021) and Popesso et al. (2023). While colibre shows good convergence with resolution at $z\gtrsim 5$, at lower redshifts the higher-resolution models tend to predict systematically lower SFRs at fixed stellar mass, leading to larger deviations from the observations. This lower normalization of the SFMS in the higher-resolution simulations appears to correlate with a slightly higher normalization of their GSMFs at similar or lower $M_{*}$ and corresponding redshifts (see Fig. 1), which may point to small but systematic differences in stellar mass growth between different resolution levels.
The fact that colibre falls short of the SFMS from Thorne et al. (2021) and Popesso et al. (2023) by 0.20.60.2-0.6 dex near cosmic noon warrants further discussion. As previously mentioned, a discrepancy of this magnitude between theoretical models and observations at $1<z<4$ is common in galaxy formation studies, and is believed to stem largely from methodological inconsistencies between theoretical predictions and observations. One way to mitigate these discrepancies is therefore to mock-observe simulated galaxies rather than relying on the internal SFRs and stellar masses predicted by the simulations. This approach involves generating realistic SEDs based on the distribution of stellar particle ages and metallicities, which are then processed through the same observational pipelines used for real data to infer SFRs and stellar masses, ensuring a more consistent comparison. By applying such a method to the eagle simulations, Katsianis et al. (2020) showed that the common tension between observed and simulated SFMS relations is largely alleviated.
@@
To sum up, in Fig. 8 we find that colibre reproduces the observed SFMS reasonably well at $0<z<1$ and $z>4$, whereas at intermediate redshifts the discrepancy between colibre and observations is $\approx 0.2-0.6$ dex. While we cannot rule out the possibility that this discrepancy is due to intrinsic limitations of the simulations, and that accounting for systematic errors in observations might even exacerbate it, the growing evidence in the literature suggests that the discrepancy between simulations and observations around cosmic noon is dominated by systematic errors on the observational side, which can likely be alleviated by moving simulated galaxies into observational space and/or by improving observational techniques to reduce these systematic errors.
Figure 9: Evolution of the galaxy quenched fraction from $z=0.1$ to $z=7$ (different panels). Thick solid lines show the quenched fraction in the simulations, where galaxies are considered quenched if their $\mathrm{sSFR}<0.2/t_{\rm H}(z)$. Shaded regions indicate the uncertainty in the predicted quenched fractions, calculated as the Clopper–Pearson interval at the 68 per cent confidence level. At low stellar masses ($M_{*}\lesssim 10^{9}~\mathrm{M_{\odot}}$), the solid lines become dotted to indicate that the increase in quenched fraction may be an artefact of limited numerical resolution. The lines also become dotted at high $M_{*}$ where the number of objects per bin drops below 10. Black symbols represent a compilation of observational data from Muzzin et al. (2013), Bauer et al. (2013), Moster et al. (2018), and Behroozi et al. (2019). Overall, colibre shows good agreement with the data at all redshifts. However, at $z\sim 0$, the most massive galaxies ($M_{*}\gtrsim 10^{11.5}~\mathrm{M_{\odot}}$) appear more star-forming than observed.
@@ 3.6 Evolution of the galaxy quenched fraction
We show the fraction of quenched galaxies as a function of stellar mass in Fig. 9. Each panel corresponds to a different redshift, ranging from $z=0.1$ (top left) to $z=7$ (bottom right). The quenched fraction in the simulations (shown as solid lines) is defined as the fraction of galaxies within a given stellar mass bin with sSFRs below $0.2/t_{\rm H}(z)$ (black dotted lines in Fig. 7). Such thresholds – which compare galaxy sSFRs to a fraction of either the Hubble time or the age of the Universe at a given redshift – are commonly used in the literature to distinguish between quiescent and star-forming galaxy populations (e.g. Franx et al., 2008; Gallazzi et al., 2014; Baker et al., 2025a; Carnall et al., 2024; Russell et al., 2025). At low stellar masses ($M_{*}\lesssim 10^{9}~\mathrm{M_{\odot}}$), the solid lines become dotted to emphasize that the rise in quenched fractions may be driven by limited numerical resolution (see below). The solid lines also become dotted at the high-mass end, indicating that there are fewer than 10 objects per stellar mass bin. The shaded regions indicate the uncertainty in the quenched fractions, computed as the Clopper–Pearson interval at the 68 per cent confidence level.
@@
To better understand the shape of the quenched fraction – $M_{*}$ relation at the high-mass end and the importance of AGN feedback in driving it, we re-ran the colibre m7 simulation in a ($50~\mathrm{cMpc}$)3 cosmological volume without SMBHs. This test confirmed that the rise in the quenched fraction from $M_{*}\sim 10^{10}$ to $\sim 10^{11.5}~\mathrm{M_{\odot}}$ is driven by AGN feedback in massive haloes, while the decrease at $M_{*}\gtrsim 10^{11.5}~\mathrm{M_{\odot}}$ is likely an indication of cooling flows in the BCGs, which the colibre AGN feedback model cannot fully counteract, resulting in overly high SFRs in these objects. However, the dip at $M_{*}\gtrsim 10^{11.5}~\mathrm{M_{\odot}}$ disappears when we increase the assumed lognormal scatter on the stellar mass from $\approx 0.1$ dex (as given by equation 1 at $z\approx 0$) to $\gtrsim 0.3$ dex (see Fig. 19 in Appendix D). Similarly, the dip also disappears if the SFR is computed within an aperture of $\lesssim 10$ pkpc (comparable to the physical size at $z=0.1$ of a fibre with a diameter of a few arcseconds in surveys such as SDSS or GAMA), rather than the default $50$ pkpc aperture used for all quantities in this work (not shown).
@@
At higher redshifts ($z\gtrsim 1$), colibre remains in reasonably good agreement with the observational data for $M_{*}\gtrsim 10^{9}~\mathrm{M_{\odot}}$, with the quenched fraction steadily decreasing with increasing redshift at all stellar masses. However, caution should be taken when drawing more detailed conclusions, as differences in the criteria for defining quiescent galaxy populations between the observations and colibre may influence the results. For galaxies with $M_{*}\sim 10^{9}-10^{10}~\mathrm{M_{\odot}}$, colibre predicts a nearly zero quenched fraction by $z=1$, which remains negligible also at higher redshifts. For $M_{*}\sim 10^{11}~\mathrm{M_{\odot}}$, the quenched fraction decreases from $\approx 70-80$ per cent at $z=0.1$ to 204020-40 per cent by $z=2$, and further drops to 5305-30 per cent by $z=5$.
Although the $z=0.1$ quenched fractions show good convergence with resolution for $10^{10}\lesssim M_{*}/\mathrm{M_{\odot}}\lesssim 10^{11}$, at $z\geq 1$ the higher-resolution simulations predict systematically lower quenched fractions (i.e. weaker quenching) than m7 in this mass range. The explanation for this discrepancy is likely twofold. First, the gas surrounding SMBHs in the higher-resolution simulations is, on average, denser than in the m7 simulation, which may increase radiative cooling losses, thereby reducing the effectiveness of AGN feedback on the host galaxy. Second, according to the colibre prescription for thermal AGN feedback, BHs inject energy into neighbouring gas particles such that the gas temperature increases by $\Delta T_{\rm AGN}$ upon injection. Because the gas particle mass is a factor of 8 (64) lower in the m6 (m5) simulation compared to m7, the energy required to heat a gas particle by the same $\Delta T_{\rm AGN}$ is also 8 (64) times lower in m6 (m5) than in m7. In other words, less energy is injected per AGN event in higher-resolution simulations, making the AGN feedback less bursty and potentially reducing its efficiency in expelling gas and quenching the galaxy. This may also explain why, at $z=0.1$, the dip in the galaxy quenched fraction at $M_{*}\gtrsim 10^{11.5}~\mathrm{M_{\odot}}$ is slightly more pronounced at m6 resolution than at m7.
@@
Satellites in colibre exhibit systematically higher quenched fractions than centrals at fixed stellar mass, a trend consistent at all resolutions. For galaxies with $M_{*}\sim 10^{9}-10^{10}~\mathrm{M_{\odot}}$, the quenched fraction of centrals ranges between 0 and $10$ per cent, whereas satellites show values of $\approx 15-40$ per cent, reflecting the important role of environmental effects. The impact of environmental effects on the quenched fraction diminishes for more massive satellites ($M_{*}\gtrsim 10^{10}~\mathrm{M_{\odot}}$). This can be explained by the deeper gravitational potential wells of more massive satellites, which allow them to better retain their gas reservoirs within the host environment (as seen in some numerical simulations and observations; e.g., Goubert et al. 2024, though see also Wetzel et al. 2012; Kukstas et al. 2023), and by the fact that they are, on average, accreted later and therefore spend less time within their host haloes. Instead, quenching in these massive satellites should primarily be driven by internal processes (in particular, AGN feedback), in a manner similar to central galaxies. As a result, the quenched fractions of massive satellites become increasingly similar to those of central galaxies at comparable stellar masses. A detailed study of environmental effects on the galaxy quenched fraction and other galaxy properties will be presented in Bahé et al. (in preparation).
For central galaxies, the quenched fraction shows good convergence in the mass range $10^{10}\lesssim M_{*}/\mathrm{M_{\odot}}\lesssim 10^{11}$ and reasonable convergence at the low ($M_{*}\lesssim 10^{10}~\mathrm{M_{\odot}}$) and high ($M_{*}\gtrsim 10^{11}~\mathrm{M_{\odot}}$) mass ends, where the higher-resolution simulations tend to predict somewhat lower values. For satellite galaxies, the convergence is worse: in the $10^{9}\lesssim M_{*}/\mathrm{M_{\odot}}\lesssim 10^{10}$ range, the quenched fraction in the m5 model is a factor of $\approx 2$ lower than in the m6 and m7 counterparts. The cause for this discrepancy is the relatively small cosmological volume used in the m5 simulation, which biases the predicted satellite quenched fraction towards lower values due to the lack of massive haloes ($M_{\rm halo}\gtrsim 10^{13.5}~\mathrm{M_{\odot}}$). In Fig. 16 of Appendix A, we show that the bias is significantly reduced when all three resolutions are compared within the same cosmological volume $(25^{3}~\mathrm{cMpc}^{3}$).